OpenStudy (anonymous):

for what values is theta is tangent 0? I really am not sure how to answer this and could really use some help

OpenStudy (anonymous):

it is the same as asking "for what values of \(x\) is \(\sin(x)=0\) as \[\tan(x)=\frac{\sin(x)}{\cos(x)}\]

OpenStudy (anonymous):

so tan =1/0 ?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

no tan = 0/1 @satellite73

OpenStudy (anonymous):

yes to the second one

OpenStudy (anonymous):

or \(x=0,\pi,2\pi,3\pi,...\)

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

@satellite73